# Exponentiations of ultrafilters

### v3

@@ 1   Introduction
A large class of problems in arithemtic Ramsey theory regards the so-called partition regular configurations and equations on =$\mathbb{N}=\{1,2,\dots\}$.
@@
Let $f_{1}\left(x_{1},\dots,x_{n}\right),\dots,f_{m}\left(x_{1},\dots,x_{n}\right):\mathbb{N}^{n}\rightarrow\mathbb{N}$. We say that the configuration $\{f_{1}\left(x_{1},\dots,x_{n}\right),\dots,f_{m}\left(x_{1},\dots,x_{n}\right)\}$ is partition regular on $\mathbb{N}$ (PR from now on) if for all finite partitions $\mathbb{N}=A_{1}\cup\dots\cup A_{k}$ there are $i\leq k$ and $x_{1},\dots,x_{n}\in\mathbb{N}$ such that for all $j\leq m$ $f_{j}\left(x_{1},\dots,x_{n}\right)\in A_{i}$.
Analogously, we say that the system $\sigma\left(x_{1},\dots,x_{n}\right)$ defined by
$\begin{cases}f_{1}\left(x_{1},\dots,x_{n}\right)=0,\\ \ \ \ \ \ \ \ \ \ \vdots\\ f_{m}\left(x_{1},\dots,x_{n}\right)=0\end{cases}$
is PR if for all finite partitions $\mathbb{N}=A_{1}\cup\dots\cup A_{k}$ there are $i\leq k$ and $x_{1},\dots,x_{n}\in A_{i}$ such that $\sigma\left(x_{1},\dots,x_{n}\right)=0$.
@@
Theorem 1.2 (Additive Schur Theorem).
@@
Equivalently, the additive Schur Theorem states that the equation $x+y=z$ is PR.
Applying transformations to partitions allows to easily deduce new results by just considering the shape of transformed partition regular configurations. A useful trick, that sometimes allows turning additive structures in multiplicative structures, and multiplicative structures in exponential ones, is that of considering $\log_{2}$-base partitions: for $B\subseteq\mathbb{N}$, let $2^{-B}=\{n\in\mathbb{N}\mid 2^{n}\in B\}$ and, given a partition $\mathbb{N}=A_{1}\cup\dots\cup A_{k}$, let the associated $\log_{2}$-base partition be given by $\mathbb{N}=2^{-A_{1}}\cup\dots\cup 2^{-A_{k}}$. For example, given any finite partition of $\mathbb{N}$, by Theorem 1.2 in the $\log_{2}$-base partition we find monochromatic $x,y,z$ with $x+y=z$; hence, $2^{x},2^{y},2^{z}=2^{x+y}=2^{x}\cdot 2^{y}$ were monochromatic with respect to the original partition. This proves the multiplicative version of Schur.
Theorem 1.3 (Multiplicative Schur Theorem).
@@
Recently, in [8], the question about the partition regularity of $\{x,y,x^{y}\}$, which we will call an exponential triple, was posed and partially answered, by using ultrafilters, by Sisto; its full solution was first given in [6] by Sahasrabudhe who, among other generalizations, proved the exponential Schur Theorem [^2] At the best of my knowledge, this theorem has not been called explicitly like this in the literature before.:
Theorem 1.4 (Exponential Schur Theorem).
@@
Again, the above theorem can be rephrased in terms of equations: it states that $x^{y}=z$ is PR. Sahasrabudhe’s proof used repeated applications of van der Waerden theorem and the compactness property on the convergence of subsequences in the space of $k$-partitions of $\mathbb{N}$. One of the key ingredients in his proof was the use of the function $f(n,m):=2^{n}m$, combined with associated $\log_{2}$-base partitions. Starting from observations about this function, in [2, Theorem 1.1] Di Nasso and Ragosta found an ultrafilters proof of the partition regularity of the configuration $\{x,y,x2^{y}\}$, from which the exponential Schur’s theorem can be immediately deduced by the same $\log_{2}$-partition trick that allows proving the multiplicative version of Schur Theorem from the additive one (see [2, Corollary 1.3]). Their proof was based on a clever trick based on tensor products of ultrafilters [^3] Through the paper we assume the reader to have a basic knowledge of the main properties of $\beta\mathbb{N}$; we refer to the monograph [5] for an expensive presentation of ultrafilters and their applications to combinatorics. and the extension to $\left(\beta\mathbb{N}\right)^{2}$ of the same function $f(n,m)=2^{n}m$ used by Sahasrabudhe is applied. Di Nasso and Ragosta used ultrafilters methods also to study the partition regularity of infinite exponential configurations in [3]. As its often the case with partition regularity problems, the ultrafilters approach provides a simple and informative proof that allows for several generalizations: this is a common situation, as proven by an enormous literature on the subject (instead of citing here hundreds of papers, we refer to [5] and its bibliography).
Our main goal is to start a systematic study of which ultrafilters are related to exponential triples, and what other combinatorial properties such ultrafilters have. This will be done by introducing two different exponentiations of ultrafilters, and by studying their properties. In particular, we will prove that all multiplicatively piecewise syndetic sets contain exponential triples which, combined with our exponentiations of ultrafilters, will allow to show that large families of ultrafilters witness the exponential Schur Theorem, in the sense that all their sets contain exponential triples.
When a binary operations on $\beta\mathbb{N}$ is introduced, it is natural to search for idempotents since, as it is well-known since Galvin and Glazer’s proof of Hindman Theorem, such idempotents have plenty of interesting combinatorial properties. Henceforth, we study and solve negatively (completely for one extension, partially for the other) the problem of the existence of idempotents in Section 2. It turns out that idempotents for such exponentiations are strictly related to idempotents for other exponential-like operations, namely generalizations of the operation $(p,q)\rightarrow 2^{p}\odot q$, which first appeared in [2].
The novelty of most of our results generates a large number of open questions and future possible directions. We conclude the paper with a discussion of many of these in Section 3.
@@ 2   Exponentiations of ultrafilters
As ultrafilters are the key notion of this paper, we assume the reader to have a basic knowledge of the algebra of $\beta\mathbb{N}$. We just briefly recall a few definitions, mostly so that we can fix a few notations.
A notion that we will use often is that of tensor product of ultrafilters in $\beta\mathbb{N}$, defined as follows:
@@
Moreover, we will need to talk about extensions of functions to $\beta\mathbb{N}$: for $f:\mathbb{N}^{n}\rightarrow\mathbb{N}$, we will denote by $\widehat{f}$ its unique continuous extension $\widehat{f}:\beta(\mathbb{N}^{n})\rightarrow\beta\mathbb{N}$. As usual, to simplify notations we identify each natural number $n\in\mathbb{N}$ with the principal idempotent it generates, so that $\mathbb{N}\subseteq\beta\mathbb{N}$.
The only purely combinatorial notion that we will need is that of multiplicatively piecewise set, defined as follows. A set $A\subseteq\mathbb{N}$ is multiplicatively syndetic if there exists a finite set $F$ such that for all $n\in\mathbb{N}$ there exists $f\in F$ such that $fx\in A$; $A$ is multiplicatively thick if for all $F$ finite subset of $\mathbb{N}$ there exists $x\in\mathbb{N}$ such that $xF\subseteq A$; $A$ is multiplicatively piecewise syndetic if $A=B\cap C$ where $B$ is multiplicatively syndetic and $C$ is multiplicatively thick. The relevant property of multiplicatively piecewise sets for us is the following:
Proposition 2.2.
[5, Theorem 4.40] $A\subseteq\mathbb{N}$ is multiplicatively piecewise syndetic if and only if there exists $p\in K(\beta\mathbb{N},\odot)$ such that $A\in p$.
@@
and the reverse sum $p\tilde{\oplus}q:=q\oplus p$; if $S(n,m):=n+m$ denotes the sum of natural numbers, then $p\oplus q=\widehat{S}(p\otimes q)$ and $q\oplus p=\widehat{S}(p\otimes q)$. Whilst different, $\oplus$ and $\tilde{\oplus}$ have the same properties, modulo properly translating their meaning by switching “left” with “right” notions. This is a consequence of the sum being commutative on $\mathbb{N}$.
@@
$E_{1}(p,q):=\{A\subseteq\mathbb{N}\mid\{n\in\mathbb{N}\mid\{m\in\mathbb{N}\mid n^{m}\in q\}\in p\},$
@@
$E_{2}(p,q):=\{A\subseteq\mathbb{N}\mid\{n\in\mathbb{N}\mid\{m\in\mathbb{N}\mid m^{n}\in q\}\in p\}.$
Equivalently, if $e:\mathbb{N}^{2}\rightarrow\mathbb{N}$ is the exponentiation $e(n,m)=n^{m}$, then $E_{1}(p,q)=\widehat{e}(p\otimes q)$, whilst $E_{2}(p,q)=\widehat{e}(q\otimes p)$.
For $n,m\in\mathbb{N}$, trivially we have that $E_{1}(n,m)=E_{2}(m,n)=n^{m}$; more in general, for $n\in\mathbb{N},p\in\beta\mathbb{N}$ we have that $E_{1}(n,p)=E_{2}(p,n):=p^{n}$ and $E_{1}(p,n)=E_{2}(n,p)=n^{p}$ as, for all $A\subseteq\mathbb{N}$
@@
Again, in terms of extension of functions, the above shows that, for $n$ fixed, $p^{n}=\widehat{e_{n}}(p)$, where $e_{n}:\mathbb{N}\rightarrow\mathbb{N}$ is the function that maps $m$ to $m^{n}$, and $n^{p}=\widehat{e^{\prime}_{n}}(p)$, where $e^{\prime}_{n}:\mathbb{N}\rightarrow\mathbb{N}$ is the function that maps $m$ to $n^{m}$.
Notation: we will continue to use the simplified notation $n^{p}$ to denote any of $E_{1}(n,p),E_{2}(p,n)$, and we will use the simplified notation $p^{n}$ to denote any of $E_{1}(p,n),E_{2}(n,p)$.
At first sight, it might look like the equality $E_{1}(p,q)=E_{2}(q,p)$ could be true for all $p,q\in\beta\mathbb{N}$; however, this is not the case. In fact, for example, let both $p,q$ contain the set $P$ of primes. Then
@@
but $P_{1}\cap P_{2}=\emptyset$, so $E_{1}(p,q)\neq E_{2}(q,p)$. We will provide some examples of nonprincipal ultrafilters $p,q$ satisfying $E_{1}(p,q)=E_{2}(q,p)$ after Proposition 2.4.
$E_{1}$ and $E_{2}$ allow to extend, in a proper way, several algebraical properties of exponentiation from natural numbers to ultrafilters. We list here those that we will need in the rest of the paper. To avoid cumbersome reformulations, some of the properties proven in the following proposition are stated for rational numbers and ultrafilters on the rationals, to which the exponentiation functions $E_{1},E_{2}$ can be extended trivially.
@@
Let $p,q,r\in\beta\mathbb{N}\setminus\{0,1\},n,m\in\mathbb{N}\setminus\{1\},a\in\mathbb{Q}\setminus\{0\}$. The following properties hold:
@@
    $n^{p}\odot n^{q}=n^{p\oplus q}$, but $E_{1}(r,p)\odot E_{1}(r,q)\neq E_{1}(r,p\oplus q)$ and $E_{2}(p,r)\odot E_{2}(q,r)\neq E_{2}(p\oplus q,r)$ in general;
@@
    $E_{2}\left(p,E_{2}(q,r)\right)=E_{2}\left(p\odot q,r\right)$; in particular, if $p$ is multiplicatively idempotent, $E_{2}\left(p,E_{2}(p,q)\right)=E_{2}(p,q)$.
@@
All these proofs consist, basically, of straightforward checks on the definitions.
$(1)$ Let $A\subseteq\mathbb{N}$. Then
$A\in E_{1}\left(p^{a},q\right)\Leftrightarrow\{n\in\mathbb{N}\mid\{m\in\mathbb{N}\mid n^{m}\in A\}\in q\}\in p^{a}\Leftrightarrow$
$\{n\in\mathbb{N}\mid n^{a}\in\{l\in\mathbb{N}\mid\{m\in\mathbb{N}\mid l^{m}\in A\}\in q\}\}\in p\Leftrightarrow\{n\in\mathbb{N}\mid\{m\in\mathbb{N}\mid\left(n^{a}\right)^{m}\in A\}\in q\}\in p\Leftrightarrow$
$\{n\in\mathbb{N}\mid\{m\in\mathbb{N}\mid\left(n^{am}\right)\in A\}\in q\}\in p\Leftrightarrow\{n\in\mathbb{N}\mid\{m\in\mathbb{N}\mid\left(n^{m}\right)\in A\}\in aq\}\in p\Leftrightarrow$
$A\in E_{1}\left(p,aq\right).$

### v4

@@ 1   Introduction
A large class of problems in arithmetic Ramsey theory regards the so-called partition regular configurations and equations on =$\mathbb{N}=\{1,2,\ldots\}$.
@@
Let $f_{1}\left(x_{1},\ldots,x_{n}\right),\ldots,f_{m}\left(x_{1},\ldots,x_{n}\right):\mathbb{N}^{n}\rightarrow\mathbb{N}$. We say that the configuration $\{f_{1}\left(x_{1},\ldots,x_{n}\right),\ldots,f_{m}\left(x_{1},\ldots,x_{n}\right)\}$ is partition regular on $\mathbb{N}$ (PR from now on) if for all finite partitions $\mathbb{N}=A_{1}\cup\ldots\cup A_{k}$ there are $i\leq k$ and $x_{1},\ldots,x_{n}\in\mathbb{N}$ such that for all $j\leq m$ $f_{j}\left(x_{1},\ldots,x_{n}\right)\in A_{i}$.
Analogously, we say that the system $\sigma\left(x_{1},\ldots,x_{n}\right)$ defined by
$\begin{cases}f_{1}\left(x_{1},\ldots,x_{n}\right)=0,\\ \ \ \ \ \ \ \ \ \ \vdots\\ f_{m}\left(x_{1},\ldots,x_{n}\right)=0\end{cases}$
is PR if for all finite partitions $\mathbb{N}=A_{1}\cup\ldots\cup A_{k}$ there are $i\leq k$ and $x_{1},\ldots,x_{n}\in A_{i}$ such that $\sigma\left(x_{1},\ldots,x_{n}\right)=0$.
@@
Theorem 1.2 (Additive Schur’s Theorem).
@@
Equivalently, the additive Schur’s Theorem states that the equation $x+y=z$ is PR.
Applying transformations to partitions allows to easily deduce new results by just considering the shape of transformed partition regular configurations. A useful trick, that sometimes allows turning additive structures in multiplicative structures, and multiplicative structures in exponential ones, is that of considering $\log_{2}$-base partitions: for $B\subseteq\mathbb{N}$, let $2^{-B}=\{n\in\mathbb{N}\mid 2^{n}\in B\}$ and, given a partition $\mathbb{N}=A_{1}\cup\ldots\cup A_{k}$, let the associated $\log_{2}$-base partition be given by $\mathbb{N}=2^{-A_{1}}\cup\ldots\cup 2^{-A_{k}}$. For example, given any finite partition of $\mathbb{N}$, in the $\log_{2}$-base partition we find monochromatic $x,y,z$ with $x+y=z$ by Theorem 1.2; hence, $2^{x},2^{y},2^{z}=2^{x+y}=2^{x}\cdot 2^{y}$ were monochromatic with respect to the original partition. This proves the multiplicative version of Schur.
Theorem 1.3 (Multiplicative Schur’s Theorem).
@@
Recently, in [9], the question about the partition regularity of $\{x,y,x^{y}\}$, which we will call an exponential triple, was posed and partially answered, by using ultrafilters, by Sisto; its full solution was first given in [7] by Sahasrabudhe who, among other generalizations, proved the exponential Schur’s Theorem [^2] At the best of our knowledge, this theorem has not been called explicitly like this in the literature before.:
Theorem 1.4 (Exponential Schur’s Theorem).
@@
Again, the above theorem can be rephrased in terms of equations: it states that $x^{y}=z$ is PR. Sahasrabudhe’s proof used the compactness property on the convergence of subsequences in the space of $k$-partitions of $\mathbb{N}$ and repeated applications of van der Waerden’s Theorem:
Theorem 1.5 (Van der Waerden’s Theorem).
For all $k\in\mathbb{N}$ the configuration $\{x,x+y,\ldots,x+ky\}$ is PR.
One of the key ingredients in Sahasrabudhe’s proof was the use of the function $f(n,m):=2^{n}m$, combined with associated $\log_{2}$-base partitions. Starting from observations about this function, in [2, Theorem 2.1] Di Nasso and Ragosta gave a very simple and elegant ultrafilters proof of the partition regularity of the configuration $\{x,y,2^{x}y\}$, from which the exponential Schur’s theorem can be immediately deduced by the same $\log_{2}$-partition trick that allows proving the multiplicative Schur’s Theorem from the additive one (see [2, Corollary 2.3]). Their proof was based on tensor products of ultrafilters [^3] We will recall some basic facts and definitions regarding ultrafilters in Section 2.1, but we assume the reader to have a basic knowledge of the main properties of $\beta\mathbb{N}$; we refer to the monograph [5] for a detailed presentation of ultrafilters and their applications to combinatorics. and the extension to $\left(\beta\mathbb{N}\right)^{2}$ of the same function $f(n,m)=2^{n}m$ used by Sahasrabudhe. Di Nasso and Ragosta used ultrafilters methods also to study the partition regularity of infinite exponential configurations in [3]. As it is often the case with partition regularity problems, the ultrafilters proof allows for several generalizations: this is a common situation, as proven by an enormous literature on the subject (instead of citing here hundreds of papers, we refer to [5] and its bibliography).
Our main goal is to start a systematic study of which ultrafilters are related to exponential triples, and what other combinatorial properties such ultrafilters have. This will be done by introducing, in Section 2, two different exponentiations of ultrafilters, named $E_{1},E_{2}$, and by studying their algebraic properties and their connections with the exponential Schur’s Theorem. In Section 2.3 we study the existence of idempotents for $E_{1},E_{2}$, and prove that $E_{1}$ does not admit nonprincipal idempotents, whilst $E_{2}$ does not admit nonprincipal idempotents that are simultaneously multiplicative idempotents. The novelty of most of our results generates a large number of open questions and future possible directions. We conclude the paper with a discussion of many of these in Section 3.
@@ 2   Exponentiations of ultrafilters
Preliminaries
As ultrafilters are the key notion of this paper, we assume the reader to have a basic knowledge of the algebra of $\beta\mathbb{N}$. We just briefly recall a few definitions, mostly so that we can fix some notations.
Sums and products of ultrafilters will be denoted by $\oplus$,$\odot$ , ⊙ respectively. Those notion are related to that of tensor product of ultrafilters in $\beta\mathbb{N}$, which is defined as follows:
@@
Moreover, we will need to talk about extensions of functions to $\beta\mathbb{N}$: for $f:\mathbb{N}^{n}\rightarrow\mathbb{N}$, we will denote by $\widehat{f}$ its unique continuous extension $\widehat{f}:\beta(\mathbb{N}^{n})\rightarrow\beta\mathbb{N}$. For example, when $S(x,y)=x+y$ is the sum, we have that $\hat{S}(p\odot q)=p\oplus q$ for all $p,q\in\beta\mathbb{N}$. As usual, to simplify notations we identify each natural number $n\in\mathbb{N}$ with the principal idempotent it generates, so that $\mathbb{N}\subseteq\beta\mathbb{N}$.
As well-known, ultrafilters are often a fundamental tool to study PR problems. For configurations, the connection between ultrafilters and partition regularity reads as follows:
Fact 2.2.
Let $f_{1}\left(x_{1},\ldots,x_{n}\right),\ldots,f_{m}\left(x_{1},\ldots,x_{n}\right):\mathbb{N}^{n}\rightarrow\mathbb{N}$. The following are equivalent:
  1. 1.
    $\{f_{1}\left(x_{1},\ldots,x_{n}\right),\ldots,f_{m}\left(x_{1},\ldots,x_{n}\right)\}$ is PR;
  2. 2.
    $\exists p\in\beta\mathbb{N}\ \forall A\in p\ \exists x_{1},\ldots,x_{n}\in\mathbb{N}\ f_{1}\left(x_{1},\ldots,x_{n}\right),\ldots,f_{m}\left(x_{1},\ldots,x_{n}\right)\in A$.
The above fact leads to the following definition:
Definition 2.3.
Let $p\in\beta\mathbb{N}$ and let $f_{1}\left(x_{1},\ldots,x_{n}\right),\ldots,f_{m}\left(x_{1},\ldots,x_{n}\right):\mathbb{N}^{n}\rightarrow\mathbb{N}$. We say that $p$ witnesses the partition regularity of the configuration $\{f_{1}\left(x_{1},\ldots,x_{n}\right),\ldots,f_{m}\left(x_{1},\ldots,x_{n}\right)\}$ if for all $A\in p$ there exist $x_{1},\ldots,x_{n}\in A$ such that $f\left(x_{1},\ldots,x_{n}\right)=0$. In this case, we write $p\models\{f_{1}\left(x_{1},\ldots,x_{n}\right),\ldots,f_{m}\left(x_{1},\ldots,x_{n}\right)\}$.
In the next sections we will use a few specific results from other papers. For the readers easiness, we recall them here. The first states that when the PR of different systems of equations are witnessed by the same ultrafilter, these configurations can be mixed.
Lemma 2.4.
[1, Lemma 2.1] Let $p\in\beta\mathbb{N}$, and let $f\left(x_{1},\ldots,x_{n}\right),g\left(y_{1},\ldots,y_{m}\right)$ be two functions such that $p\models\left(f_{1}\left(x_{1},\ldots,x_{n}\right)=0\right)\wedge\left(g\left(y_{1},\ldots,y_{m}\right)=0\right)$. Then $p\models\left(f_{1}\left(x_{1},\ldots,x_{n}\right)=0\right)\wedge\left(g\left(y_{1},\ldots,y_{m}\right)=0\right)\wedge\left(x_{1}=y_{1}\right)$.
The second result that we will need is [4, Theorem 1.6], which states that multiplicatively piecewise syndetic sets contains some exponential structure.
Theorem 2.5.
[4, Theorem 1.6] Let $n\in\mathbb{N},n>1$ and let $A\subseteq\mathbb{N}$ be multiplicatively piecewise syndetic. Then there exists $x,y\in A$ such that $xn^{y}\in A$.
The third result we need is [6, Theorem A], that states that certain ultrafilter equations admits no solution.
Theorem 2.6.
[6, Theorem A] Let $p\in\beta\mathbb{N}\setminus\mathbb{N}$, $a,b\in\mathbb{N}$ and $u,v\in\beta\mathbb{N}$. Suppose that $u\oplus ap=v\oplus bp$. Then $a=b$.
Finally, we will the non-PR of combined sums and exponentiations stated in [7, Theorem 6].
Theorem 2.7.
[7, Theorem 6] The configuration $\{x,y,x^{y},a,b,a+b\}$ is not PR.
@@
and the reverse sum $p\tilde{\oplus}q:=q\oplus p$; if $S(n,m):=n+m$ denotes the sum of natural numbers, then $p\oplus q=\widehat{S}(p\otimes q)$ and $q\oplus p=\widehat{S}(q\otimes p)$. Whilst different, $\oplus$ and $\tilde{\oplus}$ have the same properties, modulo properly translating their meaning by switching “left” with “right” notions. This is a consequence of the sum being commutative on $\mathbb{N}$.
@@
$E_{1}(p,q):=\{A\subseteq\mathbb{N}\mid\{n\in\mathbb{N}\mid\{m\in\mathbb{N}\mid n^{m}\in A\}\in q\}\in p\},$
@@
$E_{2}(p,q):=\{A\subseteq\mathbb{N}\mid\{n\in\mathbb{N}\mid\{m\in\mathbb{N}\mid m^{n}\in A\}\in q\}\in p\}.$
Equivalently, if $e_{1},e_{2}:\mathbb{N}^{2}\rightarrow\mathbb{N}$ are the exponentiations $e_{1}(n,m)=n^{m},e_{2}(n,m)=m^{n}$, then $E_{1}(p,q)=\widehat{e_{1}}(p\otimes q)$ and $E_{2}(p,q)=\widehat{e_{2}}(p\otimes q)$.
For $n,m\in\mathbb{N}$, trivially we have that $E_{1}(n,m)=E_{2}(m,n)=n^{m}$; more in general, for $n\in\mathbb{N},p\in\beta\mathbb{N}$ we have that $E_{1}(n,p)=E_{2}(p,n)$ and $E_{1}(p,n)=E_{2}(n,p)$ as, for all $A\subseteq\mathbb{N}$
@@
Notation: we will use the simplified notation $n^{p}$ to denote any of $E_{1}(n,p),E_{2}(p,n)$, and we will use the simplified notation $p^{n}$ to denote any of $E_{1}(p,n),E_{2}(n,p)$.
At first sight, it might look like the equality $E_{1}(p,q)=E_{2}(q,p)$ could be true for all $p,q\in\beta\mathbb{N}$; however, this is not the case.
Example 2.9.
Let both $p,q$ be nonprincipal ultrafilters that contain the set $P$ of primes. Then
@@
but $P_{1}\cap P_{2}=\emptyset$, so $E_{1}(p,q)\neq E_{2}(q,p)$.
We will provide some examples of nonprincipal ultrafilters $p,q$ satisfying $E_{1}(p,q)=E_{2}(q,p)$ after Proposition 2.10. To prove this proposition, it is helpful to use fractional powers of ultrafilters; to make proper sense of these powers, we prefer to state and prove Proposition 2.10 for ultrafilters in $\beta\mathbb{R}^{+}_{d}$, namely elements of the Stone-Čech compactification of +:=$\mathbb{R}^{+}:=\{x\in\mathbb{R}\mid x>0\}$ endowed with the discrete topology. We extend the exponentiations to $\beta\mathbb{R}_{d}$ as follows: given $p,q\in\beta\mathbb{R}^{+}_{d}$, we let:
$E_{1}(p,q):=\{A\subseteq\mathbb{R}^{+}\mid\{x\in\mathbb{R}^{+}\mid\{y\in\mathbb{R}^{+}\mid x^{y}\in A\}\in q\}\in p\},$
and
$E_{2}(p,q):=\{A\subseteq\mathbb{R}^{+}\mid\{x\in\mathbb{R}^{+}\mid\{y\in\mathbb{R}^{+}\mid y^{x}\in A\}\in q\}\in p\}.$
For $r\in\mathbb{R}^{+}$ and $p\in\beta\mathbb{R}^{+}_{d}$, we will keep the simplified notation $r^{p}$ to denote any of $E_{1}(r,p),E_{2}(p,r)$ (which are equal), and $p^{r}$ to denote any of $E_{1}(p,r),E_{2}(r,p)$ (which are also equal).
@@
Let $p,q,r\in\beta\mathbb{R}_{d}^{+}\setminus\{1\}$. Let $a,b\in\mathbb{R}^{+}$ with $a\geq b>1$. The following properties hold:
@@
    $a^{p}\odot a^{q}=a^{p\oplus q}$;
@@
    $E_{2}\left(p\odot q,r\right)=E_{2}\left(p,E_{2}(q,r)\right)$; in particular, if $p$ is multiplicatively idempotent, $E_{2}(p,q)=E_{2}\left(p,E_{2}(p,q)\right)$.
@@
$(1)$ Let $A\subseteq\mathbb{R}^{+}$. Then
$A\in E_{1}\left(p^{a},q\right)\Leftrightarrow\{x\in\mathbb{R}^{+}\mid\{y\in\mathbb{R}^{+}\mid x^{y}\in A\}\in q\}\in p^{a}\Leftrightarrow$
$\{z\in\mathbb{R}^{+}\mid z^{a}\in\{x\in\mathbb{R}^{+}\mid\{y\in\mathbb{R}^{+}\mid x^{y}\in A\}\in q\}\}\in p\Leftrightarrow$
$\{z\in\mathbb{R}^{+}\mid\{y\in\mathbb{R}^{+}\mid\left(z^{a}\right)^{y}\in A\}\in q\}\in p\Leftrightarrow\{z\in\mathbb{R}^{+}\mid\{y\in\mathbb{R}^{+}\mid z^{\left(ay\right)}\in A\}\in q\}\in p\Leftrightarrow$
$\{z\in\mathbb{R}^{+}\mid\{t\in\mathbb{R}^{+}\mid z^{t}\in A\}\in aq\}\in p\Leftrightarrow A\in E_{1}(p,aq).$